Simplify and express each of the following in exponential form:
 37 / (3× 33 )

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

= 37/ (34 + 3) ... [∵am × an= am + n]
= 37/ 37
= 37 – 7 ... [∵am ÷ an = am – n]
= 30
= 1

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
This can be expanded as
3/ 3+ 33
= 3/ 34+3
= 3/ 37
= 1
 
Hope it helps!
Thank you ✌️